# Xinyue Zheng Haowei Lin Kaichen He Zihao Wang Zilong Zheng Yitao Liang

### v3

@@ Abstract
Developing AI agents capable of interacting with open-world environments to solve diverse tasks is a compelling challenge. However, evaluating such open-ended agents remains difficult, with current benchmarks facing scalability limitations. To address this, we introduce Minecraft Universe (MCU), a comprehensive evaluation framework set within the open-world video game Minecraft. MCU incorporates three key components: (1) an expanding collection of 3,452 composable atomic tasks that encompasses 11 major categories and 41 subcategories of challenges; (2) a task composition mechanism capable of generating infinite diverse tasks with varying difficulty; and (3) a general evaluation framework that achieves 91.5% alignment with human ratings for open-ended task assessment. Empirical results reveal that even state-of-the-art foundation agents struggle with the increasing diversity and complexity of tasks. These findings highlight the necessity of MCU as a robust benchmark to drive progress in AI agent development within open-ended environments.
@@ 1 Introduction
Developing AI agents capable of interacting with dynamic environments—often referred to as “open-world” in the literature (Parmar et al., 2023)—to solve diverse tasks remains a long-standing challenge in Artificial Intelligence (Kejriwal et al., 2024). Among the various environments used to study AI agents, games have emerged as a prominent choice, as they provide real-world challenges within programmable simulators, offering valuable opportunities for real-world simulation (Bruce et al., 2024; Reed et al., 2022; Raad et al., 2024). Compared to other digital environments such as web/apps (Zhou et al., 2023; Qin et al., 2025; Lin et al., 2024), mobile platforms (Pan et al., 2024; Wang et al., 2024a), and coding IDEs (Xu et al., 2022; Jimenez et al., 2023; Huang et al., 2023), games present a higher degree of control complexity, similar to that encountered by physical robotic agents (Nasiriany et al., 2024; Wang et al., 2023b; Zhou et al., 2024b). While compared to robotics, games support long-horizon planning tasks and enable safer and more efficient testing within sandbox environments.
@@
Despite the numerous advantages of Minecraft as an experimental environment, we have identified several limitations in existing benchmarks that impede a comprehensive evaluation of agents’ generalizability. These limitations include low task quality (Fan et al., 2022), insufficient diversity (Hafner, 2021), and the lack of automatic evaluation suites (Cai et al., 2023). We further illustrate these issues in Figure 2. Based on these issues, we find that Minecraft agents lack a unified benchmark, and each agent is evaluated on a distinct set of tasks (see Appendix A for further discussion). To address this, we introduce Minecraft Universe (MCU), an advanced evaluation framework in Minecraft (Figure 1). MCU encompasses thousands of composable tasks and provides a scalable automatic evaluation system, designed based on the following key principles:
@@ 2.1 Minecraft Environment
    Vast State Space. Minecraft provides an extraordinarily large state space, as illustrated in Table 7 with an intuitive comparison. Its expansive maps, functional blocks, and diverse mobs result in an immense number of possible configurations. This makes Minecraft an ideal platform for studying the generalizability of agents. Additionally, we present illustrative experiments in Appendix E to demonstrate the importance of vast state spaces for agent generalization.
@@ 2.3 Atomic Tasks: Fundamental Testing Units in MCU
To pursue task diversity, as outlined in Section 1, we have collected 3,452 atomic tasks [^1] The dataset is continuously expanding., each represented as a textual description. These tasks serve as the fundamental building blocks for task synthesis in MCU. They may be later composed (discussed in the previous paragraph) and instantiated using an LLM-based task configuration generator (see Section 2.4). The atomic tasks are categorized into 11 major categories and 41 sub-categories, covering diverse challenges encountered in Minecraft. These tasks are sourced through the following pathways: (1) Distilling high-quality tasks from existing benchmarks, such as MineDojo (Fan et al., 2022) and SkillForge (Cai et al., 2023). (2) Extracting tasks from the Minecraft Wiki [^2] https://minecraft.wiki/. (3) Synthesizing tasks using in-game data from the Minecraft simulator (e.g., generating tasks like craft item X if item “X” is craftable). (4) Brainstorming innovative tasks with input from domain experts and LLMs. Further details on the collection process for atomic tasks are provided in Section C.1.
@@ 3.1 Validity of Automatic Evaluation
We collected 500 trajectories spanning 60 tasks, derived from both agent rollouts and human gameplay videos. To evaluate the quality of these trajectories, we engaged 20 expert Minecraft players to provide annotations. The players were tasked with performing both comparative evaluations and absolute ratings on randomly sampled trajectory pairs for the same task or individual trajectories. Each player contributed 1 hours of work. Details regarding the annotation process are provided in Section D.2.
Table 1: F1 scores for predicting the better human-annotated trajectory across different task categories. The compared methods include MineCLIP (Fan et al., 2022), our AutoEval without criteria (w/o criteria), and the full version of AutoEval with criteria (w/ criteria). The highest score for each task category is bolded.
@@
AutoEval (w/o criteria) 100.0 73.0 53.0 49.0 100.0 75.0
AutoEval (w/ criteria) 100.0 85.0 62.0 73.0 100.0 84.0
@@ 3.2 Evaluating Existing Agents with MCU
We evaluate four powerful foundation agents in Minecraft, all supported by MineStudio: (1) VPT (BC), a behavior cloning agent initially pre-trained on YouTube Minecraft videos and further fine-tuned on refined early-game contractor data; (2) VPT (RL), a reinforcement learning fine-tuned model based on VPT (BC) that maximizes the reward for obtaining diamonds in Minecraft; (3) STEVE-I (Lifshitz et al., 2023), which follows text instructions to complete tasks; and (4) GROOT (Cai et al., 2023), which solves tasks demonstrated by a reference video. GROOT receives video instructions, STEVE-I receives text instructions, and VPT operates without instructions. We exclude agents that simplify the Minecraft environment by deviating from the native action space. More details on these models can be found in Cai et al. (2023).
@@
While MCU enables scalable task evaluation without extensive human annotation, we carefully select a small yet diverse set of 30 atomic and 5 compositional tasks to illustrate our experimental conclusions without introducing excessive complexity. A comprehensive evaluation of a total of 150 tasks, including 90 randomly sampled atomic tasks and 60 compositional tasks, is deferred to Appendix F. The 30 atomic tasks are drawn from six major categories, ensuring inter-task diversity while maintaining a moderate difficulty level suitable for Minecraft junior players. The difficulty of each task is of simple mode, and each task is evaluated using 30 random seeds. To highlight the importance of intra-task diversity, we select six atomic tasks (1-2 from each category), assessing both simple and hard modes of these tasks seperately. Each difficulty level is evaluated with 10 random seeds.
@@ 4 Conclusion
We introduce Minecraft Universe (MCU), a scalable evaluation framework for open-ended game agents in Minecraft. MCU features 3,452 diverse and composable atomic tasks, a dynamic task composition mechanism to ensure sustained challenges, and an automated evaluation system with over 90% human alignment. Empirical results indicate that state-of-the-art agents struggle with tasks exhibiting high inter-task and intra-task diversity. By enabling scalable and high-quality evaluation, MCU establishes a robust benchmark to advance AI generalization in open-world environments. We hope MCU serves as a unified benchmark for Minecraft agent development.
@@
Appendix A Related Work
Benchmarks in Minecraft
MineDojo (Fan et al., 2022) introduces a suite of 1,560 creative tasks defined by natural language instructions. However, it suffers from significant redundancy and overly complex tasks, complicating practical evaluation, as shown in Figure 2. BEDD (Milani et al., 2023) defines five tasks covering various aspects of Minecraft, primarily designed for the MineRL BASALT competition (Shah et al., 2021). By decomposing the evaluation framework, BEDD facilitates detailed assessments of agent performance across subgoals and attributes such as human likeness. However, its reliance on human ratings limits scalability. Other works on Minecraft agents (Wang et al., 2023d, c; Cai et al., 2023; Yuan et al., 2023; Baker et al., 2022; Lifshitz et al., 2023) lack a unified benchmark, with each agent evaluated on its own task set. We argue that establishing a standardized benchmark is a high priority for advancing Minecraft agent development.
Open-ended Agents in Minecraft
Many agents have been developed to interact with Minecraft environments. Some methods focus on using Imitation Learning or Reinforcement Learning to learn various skills in open-world Minecraft to complete short-horizon tasks (Baker et al., 2022; Cai et al., 2023; Lifshitz et al., 2023; Jiang et al., 2025; Zhao et al., 2024a; Cai et al., ; Yuan et al., 2024; Cai et al., 2024b; Jiang & Lu, 2025). Baker et al. (2022) generates action labels from Minecraft videos on YouTube using IDM and performs imitation learning to obtain an unconditional policy capable of completing various tasks. Lifshitz et al. (2023) uses MineCLIP as a text encoder to obtain a text-conditioned multitask policy based on unconditional VPT (Baker et al., 2022). Minecraft also supports the testing of long-horizon programmatic tasks, so some methods accomplish long-horizon tasks by using large language models as planners, combined with skill policies (Wang et al., 2023d; Zhou et al., 2024a; Chen & Gao, 2024; Yuan et al., 2023; Qin et al., 2024; Zheng et al., 2023b). Wang et al. (2023d) has designed an agent pipeline based on GPT, which interactively completes tasks from environmental feedback through the self-explanation and zero-shot planning capabilities of LLMs. In order to further enhance the long-horizon capabilities of LLM Agents, some methods have explored efficient explicit memory mechanisms to support agents retrieve and improve from previous interaction trajectories (Wang et al., 2023c; Zhu et al., 2023; Wang et al., 2023a; Park et al., 2024; Li et al., 2024). Unlike designing complex agent pipelines, the rise of VLA (Brohan et al., 2023; Kim et al., 2024) has inspired researchers to use end-to-end VLM as a policy to directly fulfill human instructions in Minecraft (Wang et al., 2024b; Zhao et al., 2024b). Some methods use a code-as-policy approach (Liang et al., 2022), using MineFlayer (PrismarineJS, 2024) as a language-conditioned policy, combined with a pretrained LLM to accomplish various tasks to avoid agents trapped on short-horizon skills lacking when executing long-horizon tasks (Wang et al., 2023a; Yang et al., 2025; Liu et al., 2024b, a). Finally, there are some methods focused on completing open-ended creative tasks in Minecraft, such as building and decoration, which differ from the traditional programmatic object-centric tasks and often cannot be directly defined by rule-based rewards (Guo et al., 2024; Zhang et al., 2023).
LLM-as-a-Judge
Appendix B Minecraft Environment Setting
@@
B.1 Minecraft Game World Setting
@@ B.3 Action Space
B.4 Why Minecraft is Suitable for Open-Ended Agent?
@@ 36 Slots, 500+ Item Types, Max Stack 64
B.4.1 State Space Calculation
@@ B.4.2 Further Attributes
Appendix C Details of Task Generation
C.1 The Source of Atomic Tasks
@@
We curated tasks from established benchmarks, including Skill-Forge (Cai et al., 2023), MineDojo (Fan et al., 2022), and prior Minecraft research (Baker et al., 2022; Yuan et al., 2023; Wang et al., 2023d, c). To refine the selection, we performed deduplication to remove redundant tasks, excluded those too difficult for human players (e.g., constructing highly complex architectures or automated redstone circuits), and eliminated compositional tasks that could be decomposed into two or more atomic tasks.
@@ C.2.2 Inventories
Appendix D Human Annotation
D.1 Minecraft Quiz
@@
D.2 Human Videos For Tasks
@@
D.3 Human rating system
@@
Appendix E Generalization Experiments
@@ E.3 Discussion
Appendix F Large-Scale Inference and Evaluation of MCU
@@
F.1 Atomic Tasks Setup
@@
F.2 Composite Tasks Setup

### v4

# Xinyue Zheng Haowei Lin Kaichen He Zihao Wang Qiang Fu Haobo Fu Zilong Zheng Yitao Liang

@@ Abstract
Developing AI agents capable of interacting with open-world environments to solve diverse tasks is a compelling challenge. However, evaluating such open-ended agents remains difficult, with current benchmarks facing scalability limitations. To address this, we introduce Minecraft Universe (MCU), a comprehensive evaluation framework set within the open-world video game Minecraft. MCU incorporates three key components: (1) an expanding collection of 3,452 composable atomic tasks that encompasses 11 major categories and 41 subcategories of challenges; (2) a task composition mechanism capable of generating infinite diverse tasks with varying difficulty; and (3) a general evaluation framework that achieves 91.5% alignment with human ratings for open-ended task assessment. Empirical results reveal that even state-of-the-art foundation agents struggle with the increasing diversity and complexity of tasks. These findings highlight the necessity of MCU as a robust benchmark to drive progress in AI agent development within open-ended environments. Our evaluation code and scripts are available at https://github.com/CraftJarvis/MCU.
@@ 1 Introduction
Developing AI agents capable of interacting with dynamic environments—often referred to as “open-world” in the literature (Parmar et al., 2023)—to solve diverse tasks remains a long-standing challenge in Artificial Intelligence (Kejriwal et al., 2024). Among the various environments used to study AI agents, games have emerged as a prominent choice, as they provide real-world challenges within programmable simulators, offering valuable opportunities for real-world simulation (Bruce et al., 2024; Reed et al., 2022; Raad et al., 2024). Compared to other digital environments such as web/apps (Zhou et al., 2023; Qin et al., 2025; Lin et al., 2024), mobile platforms (Pan et al., 2024; Wang et al., 2024), and coding IDEs (Xu et al., 2022; Jimenez et al., 2024; Huang et al., 2023), games present a higher degree of control complexity, similar to that encountered by physical robotic agents (Nasiriany et al., 2024; Wang et al., 2023b; Zhou et al., 2024b). While compared to robotics, games support long-horizon planning tasks and enable safer and more efficient testing within sandbox environments.
@@
Despite the numerous advantages of Minecraft as an experimental environment, we have identified several limitations in existing benchmarks that impede a comprehensive evaluation of agents’ generalizability. These limitations include low task quality (Fan et al., 2022), insufficient diversity (Hafner, 2021), and the lack of automatic evaluation suites (Cai et al., 2024c). We further illustrate these issues in Figure 2. Based on these issues, we find that Minecraft agents lack a unified benchmark, and each agent is evaluated on a distinct set of tasks (see Section 4 for further discussion). To address this, we introduce Minecraft Universe (MCU), an advanced evaluation framework in Minecraft (Figure 1). MCU encompasses thousands of composable tasks and provides a scalable automatic evaluation system, designed based on the following key principles:
@@ 2.1 Minecraft Environment
    Vast State Space. Minecraft provides an extraordinarily large state space, as illustrated in Table 7 with an intuitive comparison. Its expansive maps, functional blocks, and diverse mobs result in an immense number of possible configurations. This makes Minecraft an ideal platform for studying the generalizability of agents. Additionally, we present illustrative experiments in Appendix D to demonstrate the importance of vast state spaces for agent generalization.
@@ 2.3 Atomic Tasks: Fundamental Testing Units in MCU
To pursue task diversity, as outlined in Section 1, we have collected 3,452 atomic tasks [^1] The dataset is continuously expanding., each represented as a textual description. These tasks serve as the fundamental building blocks for task synthesis in MCU. They may be later composed (discussed in the previous paragraph) and instantiated using an LLM-based task configuration generator (see Section 2.4). The atomic tasks are categorized into 11 major categories and 41 sub-categories, covering diverse challenges encountered in Minecraft. These tasks are sourced through the following pathways: (1) Distilling high-quality tasks from existing benchmarks, such as MineDojo (Fan et al., 2022) and SkillForge (Cai et al., 2024c). (2) Extracting tasks from the Minecraft Wiki [^2] https://minecraft.wiki/. (3) Synthesizing tasks using in-game data from the Minecraft simulator (e.g., generating tasks like craft item X if item “X” is craftable). (4) Brainstorming innovative tasks with input from domain experts and LLMs. Further details on the collection process for atomic tasks are provided in Section B.1.
@@ 3.1 Validity of Automatic Evaluation
We collected 500 trajectories spanning 60 tasks, derived from both agent rollouts and human gameplay videos. To evaluate the quality of these trajectories, we engaged 20 expert Minecraft players to provide annotations. The players were tasked with performing both comparative evaluations and absolute ratings on randomly sampled trajectory pairs for the same task or individual trajectories. Each player contributed 1 hours of work. Details regarding the annotation process are provided in Section C.2.
Table 1: F1 scores for predicting the better human-annotated trajectory across different task categories. The compared methods include MineCLIP (Fan et al., 2022), our AutoEval on both open-access models (MiniCPM (Yao et al., 2024), JarvisVLA (Li et al., 2025)) and closed API-based models (GPT-4o). The highest score for each task category is bolded.
@@
AutoEval (MiniCPM) 65.0 43.0 80.0 59.0 53.0 60.0
AutoEval (JarvisVLA) 73.0 62.0 73.0 84.0 65.0 71.4
AutoEval (GPT-4o) 100.0 85.0 62.0 73.0 100.0 84.0
@@ 3.2 Evaluating Existing Agents with MCU
We evaluate four powerful foundation agents in Minecraft, all supported by MineStudio: (1) VPT (BC), a behavior cloning agent initially pre-trained on YouTube Minecraft videos and further fine-tuned on refined early-game contractor data; (2) VPT (RL), a reinforcement learning fine-tuned model based on VPT (BC) that maximizes the reward for obtaining diamonds in Minecraft; (3) STEVE-I (Lifshitz et al., 2023), which follows text instructions to complete tasks; and (4) GROOT (Cai et al., 2024c), which solves tasks demonstrated by a reference video. GROOT receives video instructions, STEVE-I receives text instructions, and VPT operates without instructions. We exclude agents that simplify the Minecraft environment by deviating from the native action space. More details on these models can be found in Cai et al. (2024c).
@@
While MCU enables scalable task evaluation without extensive human annotation, we carefully select a small yet diverse set of 30 atomic and 5 compositional tasks to illustrate our experimental conclusions without introducing excessive complexity. A comprehensive evaluation of a total of 150 tasks, including 90 randomly sampled atomic tasks and 60 compositional tasks, is deferred to Appendix E. The 30 atomic tasks are drawn from six major categories, ensuring inter-task diversity while maintaining a moderate difficulty level suitable for Minecraft junior players. The difficulty of each task is of simple mode, and each task is evaluated using 30 random seeds. To highlight the importance of intra-task diversity, we select six atomic tasks (1-2 from each category), assessing both simple and hard modes of these tasks seperately. Each difficulty level is evaluated with 10 random seeds.
@@ 3.2.2 Intra-task Generalization
4 Related Work
Benchmarks in Minecraft
MineDojo (Fan et al., 2022) introduces a suite of 1,560 creative tasks defined by natural language instructions. However, it suffers from significant redundancy and overly complex tasks, complicating practical evaluation, as shown in Figure 2. BEDD (Milani et al., 2023) defines five tasks covering various aspects of Minecraft, primarily designed for the MineRL BASALT competition. By decomposing the evaluation framework, BEDD facilitates detailed assessments of agent performance across subgoals and attributes such as human likeness. However, its reliance on human ratings limits scalability. Other works on Minecraft agents (Wang et al., 2023d, c; Cai et al., 2024c; Yuan et al., 2023; Baker et al., 2022; Lifshitz et al., 2023) lack a unified benchmark, with each agent evaluated on its own task set. We argue that establishing a standardized benchmark is a high priority for advancing Minecraft agent development.
Open-ended Agents in Minecraft
Many agents have been developed to interact with Minecraft environments. Some methods focus on using Imitation Learning or Reinforcement Learning to learn various skills in open-world Minecraft to complete short-horizon tasks (Baker et al., 2022; Cai et al., 2024c; Lifshitz et al., 2023; Jiang et al., 2025; Zhao et al., 2024a; Cai et al., ; Yuan et al., 2024; Cai et al., 2024b; Jiang & Lu, 2025). Baker et al. (2022) generates action labels from Minecraft videos on YouTube using IDM and performs imitation learning to obtain an unconditional policy capable of completing various tasks. Lifshitz et al. (2023) uses MineCLIP as a text encoder to obtain a text-conditioned multitask policy based on unconditional VPT (Baker et al., 2022). Minecraft also supports the testing of long-horizon programmatic tasks, so some methods accomplish long-horizon tasks by using large language models as planners, combined with skill policies (Wang et al., 2023d; Zhou et al., 2024a; Chen & Gao, 2024; Yuan et al., 2023; Qin et al., 2024; Zheng et al., 2023b). Wang et al. (2023d) has designed an agent pipeline based on GPT, which interactively completes tasks from environmental feedback through the self-explanation and zero-shot planning capabilities of LLMs. In order to further enhance the long-horizon capabilities of LLM Agents, some methods have explored efficient explicit memory mechanisms to support agents retrieve and improve from previous interaction trajectories (Wang et al., 2023c; Zhu et al., 2023; Wang et al., 2023a; Park et al., 2024; Li et al., 2024). Unlike designing complex agent pipelines, the rise of VLA (Zitkovich et al., 2023; Kim et al., 2024) has inspired researchers to use end-to-end VLM as a policy to directly fulfill human instructions in Minecraft (omnijarvis; Zhao et al., 2024b). Some methods use a code-as-policy approach (Liang et al., 2022), using MineFlayer (PrismarineJS, 2024) as a language-conditioned policy, combined with a pretrained LLM to accomplish various tasks to avoid agents trapped on short-horizon skills lacking when executing long-horizon tasks (Wang et al., 2023a; Yang et al., 2025; Liu et al., 2024b, a). Finally, there are some methods focused on completing open-ended creative tasks in Minecraft, such as building and decoration, which differ from the traditional programmatic object-centric tasks and often cannot be directly defined by rule-based rewards (Guo et al., 2024; Zhang et al., 2023).
LLM-as-a-Judge
@@ 5 Conclusion
We introduce Minecraft Universe (MCU), a scalable evaluation framework for open-ended game agents in Minecraft. MCU features 3,452 diverse and composable atomic tasks, a dynamic task composition mechanism to ensure sustained challenges, and an automated evaluation system with over 90% human alignment. Empirical results indicate that state-of-the-art agents struggle with tasks exhibiting high inter-task and intra-task diversity. To support standardized benchmarking, we release MCU-Turbo, a curated subset of 100 tasks with structured difficulty settings, as detailed in Appendix F. We hope MCU serve as robust foundations for advancing open-world agent research.
Acknowledgement
This work was funded in part by the National Science and Technology Major Project (2022ZD0114902), and National Natural Science Foundation of China (62376031). We acknowledge other members in CraftJarvis team for their data labeling support and constructive discussions, and thank Dr. Wenzheng Feng for his valuable insights.
@@ Acknowledgement
Appendix A Minecraft Environment Setting
@@
A.1 Minecraft Game World Setting
@@ A.3 Action Space
A.4 Why Minecraft is Suitable for Open-Ended Agent?
@@ 36 Slots, 500+ Item Types, Max Stack 64
A.4.1 State Space Calculation
@@ A.4.2 Further Attributes
Appendix B Details of Task Generation
B.1 The Source of Atomic Tasks
@@
We curated tasks from established benchmarks, including Skill-Forge (Cai et al., 2024c), MineDojo (Fan et al., 2022), and prior Minecraft research (Baker et al., 2022; Yuan et al., 2023; Wang et al., 2023d, c). To refine the selection, we performed deduplication to remove redundant tasks, excluded those too difficult for human players (e.g., constructing highly complex architectures or automated redstone circuits), and eliminated compositional tasks that could be decomposed into two or more atomic tasks.
@@ B.2.2 Inventories
Appendix C Human Annotation
C.1 Minecraft Quiz
@@
C.2 Human Videos For Tasks
@@
C.3 Human rating system
@@
Appendix D Generalization Experiments
@@ D.3 Discussion
Appendix E Large-Scale Inference and Evaluation of MCU
@@
E.1 Atomic Tasks Setup
@@
E.2 Composite Tasks Setup
@@ E.3 Experimental Conclusions
Appendix F MCU-Turbo: A Standard Benchmark for Evaluating Minecraft Agents
We introduce MCU-Turbo, a canonical benchmark suite for systematically evaluating agents within the Minecraft Universal (MCU) framework. MCU-Turbo is designed as a standardized evaluation protocol, comprising 80 atomic tasks across 10 categories and 20 compositional tasks. Each task is assessed under two difficulty regimes—Simple and Hard—to rigorously test an agent’s capabilities in generalization, tool use, long-horizon planning, and robustness to environmental variation. The following experiments establish baseline performance using current state-of-the-art agents.
F.1 Baseline Results in the Simple Evaluation Mode
Agents demonstrate varied competencies across different evaluation dimensions under the simple setting. Notably, aspects such as creative behavior and error recognition remain challenging across the board, as illustrated in Table 10.
Table 10: Baseline agent performance in simple mode.
Agent Name Task Progress Action Control Error Recognition Creative Attempts Task Efficiency Material Usage
STEVE-1 31.4% 31.9% 13.1% 6.4% 23.2% 35.6%
VPT (BC) 29.1% 29.0% 11.8% 6.2% 21.3% 33.8%
VPT (RL) 25.9% 26.2% 8.8% 5.2% 18.9% 31.3%
JARVIS-VLA 25.6% 27.8% 9.3% 5.5% 18.3% 30.5%
F.2 Baseline Results in the Hard Evaluation Mode
In the Hard mode, agents are subjected to increased environmental complexity and the presence of distractors. As shown in Table 11, performance consistently declines across all dimensions, underscoring the difficulty of generalization under more challenging conditions.
Table 11: Performance degradation in hard mode.
Agent Name Task Progress Action Control Error Recognition Creative Attempts Task Efficiency Material Usage
STEVE-1 23.1% 22.6% 6.9% 6.0% 16.6% 24.5%
VPT (BC) 23.0% 21.7% 6.2% 6.0% 15.6% 25.0%
VPT (RL) 21.0% 20.9% 5.0% 4.6% 14.4% 23.7%
JARVIS-VLA 20.9% 22.3% 5.5% 4.3% 14.5% 22.3%
F.3 Agent Performance on Creative vs. Programmatic Tasks
Creative tasks present a substantially greater challenge compared to programmatic ones. For example, STEVE-1 exhibits a 15.8% reduction in task progress when transitioning from programmatic to creative tasks, as shown in Table 12. These results highlight the persistent difficulty of generalization in open-ended and less-structured settings.
Table 12: Performance comparison: creative vs. programmatic tasks.
Task Type Task Progress Action Control Error Recognition Creative Attempts Task Efficiency Material Usage
Programmatic 38.4% 36.5% 16.3% 7.7% 27.7% 43.8%
Creative 22.6% 26.2% 9.1% 4.7% 17.7% 25.4%
Drop -15.8% -10.3% -7.2% -3.0% -10.0% -18.4%
Overall, the MCU-Turbo benchmark provides fine-grained insights into agent capabilities across a diverse task spectrum. It emphasizes enduring challenges in creativity and adaptive behavior, and aims to steer future research toward the development of more general-purpose, robust Minecraft agents.